OpenStudy (isabel☺):

-(1/2)ln2 + (1/2)ln3 without using calculator ☺

OpenStudy (isabel☺):

@UnkleRhaukus @Hero

OpenStudy (isabel☺):

@ganeshie8

OpenStudy (isabel☺):

@Nursi

OpenStudy (dumbcow):

combine into 1 log \[\ln (2^{-1/2} *3^{1/2})\]

OpenStudy (isabel☺):

that's the final answer?

OpenStudy (dumbcow):

yeah i guess if they want an exact answer you can write it as \[\ln (\sqrt{\frac{3}{2}})\]

OpenStudy (isabel☺):